### v1

@@ Abstract
Policy gradient (PG) methods are the backbone of many reinforcement learning algorithms due to their good performance in policy optimization problems. As a gradient-based approach, PG methods typically rely on knowledge of the system dynamics. However, this information is not always available, and in such cases, trajectory data can be utilized to approximate first-order information. When the data are noisy, gradient estimates become inaccurate and a formal investigation that encompasses uncertainty estimation and the analysis of its propagation through the algorithm is currently missing. To address this, our work focuses on the Linear Quadratic Regulator (LQR) problem for systems subject to additive stochastic noise. After briefly summarizing the state of the art for cases with a known model, we focus on scenarios where the system dynamics are unknown, and approximate gradient information is obtained using zeroth-order optimization techniques. We analyze the theoretical properties by computing the error in the estimated gradient and examining how this error affects the convergence of PG algorithms. Additionally, we provide global convergence guarantees for various versions of PG methods, including those employing adaptive step sizes and variance reduction techniques, which help increase the convergence rate and reduce sample complexity. One contribution of this work is the study of the robustness of model-free PG methods, aiming to identify their limitations in the presence of noise and propose improvements to enhance their applicability. Numerical simulations show that these theoretical analyses provide valuable guidance in tuning the algorithm parameters, thereby making these methods more reliable in practically relevant scenarios.
keywords:
Policy Gradient Methods, Sample Complexity, Linear Optimal Control, Zeroth-order Optimization
{MSCcodes}
93E20, 90C56, 49N10
@@ 1 Introduction
Reinforcement learning [3, 4, 27] has demonstrated a significant influence across a wide range of applications. A key concept within reinforcement learning is policy optimization, where the policy is parameterized and directly optimized over based on a predefined performance metric [14]. Several successful policy optimization methods have been developed, including policy gradient [28], actor-critic [16], proximal policy optimization [23]. This work focuses on policy gradient (PG) methods, which are based on the simple idea of minimizing a cost function over the parameterized policy by improving performance through a gradient descent-type update. Studying the convergence properties of PG methods, particularly their global convergence to the optimal policy, is an active area of research [6, 10, 11, 34, 35, 36, 37]. For instance, PG methods have been applied in [34] to solve $\mathcal{H}_{2}$ cost function subject to $\mathcal{H}_{\infty}$ constraints, and they have been applied in [35] to the Markov decision processes.
@@
The application of reinforcement learning to the linear quadratic regulator (LQR) problem has been extensively explored due to its value as an analytically tractable benchmark, making it ideal for studying reinforcement learning in environments with continuous state and action spaces [2, 8, 10, 17, 20, 22, 30]. Policy gradient methods have gained attention in recent studies focusing on the LQR problem. Interest in applying these methods to the LQR setting increased significantly after the work in [10] demonstrated global convergence properties of policy gradient methods applied to the deterministic LQR problem. In [37], the performance limitations of PG applied to LQR problem are studied from a control-theoretic perspective. The work [36] explores a primal-dual policy gradient method to solve the constrained LQR problem.
@@
Solving the LQR problem requires the availability of a model of the system, as it is the case in [34, 35, 36, 37], where the analysis assumes knowledge of the exact model. However, in real-world applications, a complete system description is often unavailable, making it necessary to combine policy optimization schemes with data-driven techniques. To address this challenge, various approaches have been developed. For instance, methods like those in [1, 5, 25] combine system identification with model-based LQR design and use regret as a performance metric of the learning process. It is also possible to use the same two-step approach to solve the LQR problem with gradient-based schemes. The work [26] integrates recursive least squares with policy iteration optimization, while [24] combines recursive least squares with the policy gradient method. These approaches are termed indirect data-driven methods, because first data is used to estimate a model, which is then integrated into a model-based certainty equivalence design. An alternative category is direct data-driven control, which directly uses data to design the controller bypassing the intermediate model estimation process. For instance, a model-free policy iteration algorithm is proposed in [33]. Several model-free policy gradient methods have also been introduced and theoretically analyzed. In [10], model-free PG methods are proposed and formulated using zeroth-order optimization to estimate gradients for the LQR cost from finite but noise-free system trajectories. Similarly, in [29], PG methods are applied to a continuous-time output feedback system using zeroth-order gradient estimation alongside variance reduction techniques. In the aforementioned model-free works [10, 24, 26, 29, 33], only the case where the system’s trajectories are noise-free is studied. However, it is essential to analyze what happens when gradients and other objects involved in the policy update are estimated using noisy data. In [11], Model-free PG methods are applied to the linear system with multiplicative noise, addressing uncertainty in the system matrix. To date, there is no holistic analysis including the uncertainty associated with the estimation process and the analysis of its propagation through the algorithm, including probabilistic guarantees on convergence and its sample complexity
@@
In this work, we analyze the sample complexity and robustness to noise of PG methods for the model-free LQR with additive stochastic noise. Due to the presence of noise with unbounded support, we consider an average infinite-horizon cost. We first characterize the properties of this cost function and analyze model-based versions of a few PG algorithms. For the gradient descent (GD) and natural policy gradient (NPG) methods we propose an adaptive stepsize scheme that has a provable faster convergence rate compared to the fixed step size methods discussed in [10, 11, 29]. Then, for both the model-free gradient descent and natural policy gradient methods, we employ zeroth-order optimization to approximate the gradient and covariance from noisy data. We provide convergence guarantees for these model-free PG methods with a sample complexity that takes into account the noisy source of estimation error. Furthermore, we introduce and analyze a variance reduction technique within the zeroth-order optimization framework, and prove that this achieves a guaranteed improvement in the sample complexity. Even though variance reduction techniques for model-free LQR were already presented in [29], the contribution here represents a significant extension as we work in the noisy scenario and we are able to provide a provable improvement in the number of required samples. Finally, we provide a comparative analysis of model-free PG methods with and without noisy data, to demonstrate the importance of the analysis considering noise, offer guidance of parameter tuning in applying policy gradient methods to applications, and also illustrate the improvement of introducing adaptive step sizes and variance reduction. Our analysis shows that model-free PG with stochastic data enjoy qualitatively similar convergence guarantees to those in the noise-free scenario, but it points out limitations in terms of the number of samples necessaries, stepsize ranges and convergence rate that are caused by noise.
@@
[…]
@@ 2 Preliminaries
(1a) $\min_{u_{t}}$ $\lim_{T\rightarrow+\infty}\frac{1}{T}\mathop{\mathbb{E}}_{x_{0},w_{t}}\sum_{t=0}^{T-1}\left(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}\right),$
(1b) $\mathrm{s.\,t.}~{}$ $x_{t+1}=Ax_{t}+Bu_{t}+w_{t},x_{0}\sim\mathcal{D},$
where $x_{t}\in\mathbb{R}^{n_{x}}$ is the system state and $u_{t}\in\mathbb{R}^{n_{u}}$ is the system input; $A\in\mathbb{R}^{n_{x}\times n_{x}}$, $B\in\mathbb{R}^{n_{x}\times n_{u}}$, $(A,B)$ is unknown but stabilizable, which is a standard assumption in data-driven control [31, 33]; $w_{t}\sim\mathcal{N}(0,\Sigma_{w})$; $Q,R\succ 0$ are the weighting matrix. We define the state covariance at the initial time as $\Sigma_{0}:=\mathop{\mathbb{E}}\limits_{x_{0}}\left[x_{0}x_{0}^{\top}\right]$ and then $\mathcal{D}:=\mathcal{N}(0,\Sigma_{0})$.
@@
(2) $C(K):=\lim_{T\rightarrow+\infty}\frac{1}{T}\mathop{\mathbb{E}}_{x_{0},w_{t}}\sum_{t=0}^{T-1}x_{t}^{\top}\underbrace{\left(Q+K^{\top}RK\right)}_{=:Q_{K}}x_{t}.$
@@ 2.1 Properties of $C$
(4) $\mathcal{S}:=\left\{K\in\mathbb{R}^{n_{x}\times n_{u}}|\lVert A_{K}\rVert<1\right\}.$
@@
(5) $\Sigma_{t+1}=A_{K}\Sigma_{t}A_{K}^{\top}+\Sigma_{w},\quad t\in\mathbb{Z}_{+}.\$
@@
(6) $\Sigma_{K}:=\lim_{T\rightarrow+\infty}\frac{1}{T}\sum_{t=0}^{T-1}\Sigma_{t}.$
@@
(7) $\Sigma_{K}=\Sigma_{w}+A_{K}\Sigma_{K}A_{K}^{\top}=\sum_{t=0}^{\infty}A_{K}^{t}\Sigma_{w}A_{K}^{t~{}\top}.$
@@
(8) $C(K)=\mathrm{Tr}\left(P_{K}\Sigma_{w}\right)=\mathrm{Tr}\left(Q_{K}\Sigma_{K}\right),$
@@
It is known that the cost function $C$ defined in (2) is generally non-convex. In the context of non-convex optimization, the convergence to the optimal solution of gradient descent cannot usually be guaranteed. However, the cost function $C(K)$ satisfies a special condition known as gradient domination. {Lemma}[Gradient Domination] The function $C$ on the set $\mathcal{S}$ is gradient dominated. That is, for any $K\in\mathcal{S}$, the following inequality holds:
(12) $C(K)-C(K^{*})\leq\mu\lVert\nabla C(K)\rVert_{F}^{2},$
with $\mu:=\frac{1}{4}\lVert\Sigma_{K^{*}}\rVert\lVert\Sigma_{w}^{-2}\rVert\lVert R^{-1}\rVert$. From (12) and the fact that $\Sigma_{w}\succ 0$, it follows that the cost function $C$ has a unique minimizer and no other stationary points. The proof of Lemma 2.1 is provided in Appendix A.1, and is a straightforward extension of [10, Lemma 3] to the averaged infinite horizon setting considered here (2). {Lemma}[Almost Smoothness on $\mathcal{S}$] For any $K,K^{\prime}\in\mathbb{S}$, the following inequality holds:
(13) $\begin{split}\left\lvert C(K^{\prime})-C(K)-2\mathrm{Tr}\left(({K^{\prime}}-{K})^{\top}E_{K}\Sigma_{K^{\prime}}\right)\right\rvert\leq\lVert\Sigma_{K^{\prime}}\rVert\left\lVert R+B^{\top}P_{K}B\right\rVert\left\lVert K^{\prime}-{K}\right\rVert_{F}^{2},\end{split}$
The proof of Lemma 12 is provided in Appendix A.2, which is an extension of [10, Lemma 6]. To understand why (13) is referred to as almost smoothness, assume additionally that: we can approximate $\Sigma_{K}$ with $\Sigma_{K^{\prime}}$, i.e. $\Sigma_{K}\approx\Sigma_{K^{\prime}}$, and that we can upper bound the term $\lVert\Sigma_{K}\rVert\left\lVert R+B^{\top}P_{K}B\right\rVert$ by $L$. Then, recalling (13), we see that (11) is equivalent to:
@@ 2.2 Perturbation of $\Sigma_{K},C,\nabla C$
In this subsection, we demonstrate that the average covariance matrix $\Sigma_{K}$, the cost function $C(K)$, and its gradient $\nabla C(K)$ are locally Lipschitz continuous with respect to the policy $K$. This property is crucial for the analysis of model-free policy optimization. {Lemma}[$\Sigma_{K}$ Perturbation] Suppose $K^{\prime},K\in\mathcal{S}$ are such that:
@@
The proof of Lemma 2.2 is given in Appendix A.3, which follows the proof of [10, Lemma 16] and uses the definition of $\Sigma_{K}$.
{Lemma}
[$C$ Perturbation] Suppose $K^{\prime},K\in\mathcal{S}$ are such that:
@@
where $h_{C}(C(K))$ is defined in (55) in the proof. {Lemma}[$\nabla C$ Perturbation] Suppose $K^{\prime}$ is such that:
@@
where $h_{\nabla}(C(K))$ is defined in (59) in the proof. The proofs of Lemma 2.2 and Lemma 2.2 are provided in Appendix A.4 and A.5, respectively. These proofs follow the approach used in [10, Lemma 24, Lemma 25] by adapting the expressions of $C$ and $\nabla C$. In summary, Lemma 2.2, Lemma 2.2 and Lemma 2.2 establish that $\Sigma_{K}$, $C$ and $\nabla C$ are locally Lipschitz continuous.

### v2

@@ Abstract
Policy gradient (PG) methods are the backbone of many reinforcement learning algorithms due to their good performance in policy optimization problems. As a gradient-based approach, PG methods typically rely on knowledge of the system dynamics. If this is not available, trajectory data can be utilized to approximate first-order information. When the data are noisy, gradient estimates become inaccurate and a study that investigates uncertainty estimation and the analysis of its propagation through the algorithm is currently missing. To address this, our work focuses on the Linear Quadratic Regulator (LQR) problem for systems subject to additive stochastic noise. After briefly summarizing the state of the art for cases with a known model, we focus on scenarios where the system dynamics are unknown, and approximate gradient information is obtained using zeroth-order optimization techniques. We analyze the theoretical properties by computing the error in the estimated gradient and examining how this error affects the convergence of PG algorithms. Additionally, we provide global convergence guarantees for various versions of PG methods, including those employing adaptive step sizes and variance reduction techniques, which help increase the convergence rate and reduce sample complexity. This study contributed to characterizing robustness of the study of the robustness of model-free PG methods, aiming to identify their limitations in the presence of stochastic noise and proposing improvements to enhance their applicability.
Keywords Policy Gradient Methods $\cdot$ Sample Complexity $\cdot$ Linear Optimal Control $\cdot$ Zeroth-order Optimization
@@ 1 Introduction
Reinforcement learning [1, 2, 3] has demonstrated a significant influence across a wide range of applications. A key concept within reinforcement learning is policy optimization, whereby the policy is parameterized and directly optimized over based on a predefined performance metric [4]. Several successful policy optimization methods have been developed, including policy gradient [5], actor-critic [6], proximal policy optimization [7]. This work focuses on policy gradient (PG) methods, which are based on the simple idea of minimizing a cost function over the parameterized policy by improving performance through a gradient descent-type update. Studying the convergence properties of PG methods, particularly their global convergence to the optimal policy, is an active area of research [8, 9, 10, 11, 12, 13, 14, 15]. For instance, PG methods have been applied in [11] to solve $\mathcal{H}_{2}$ cost function subject to $\mathcal{H}_{\infty}$ constraints, and they have been applied in [12] to Markov decision processes.
@@
The application of reinforcement learning to the linear quadratic regulator (LQR) problem has been extensively explored due to its value as an analytically tractable benchmark, making it ideal for studying reinforcement learning in environments with continuous state and action spaces [16, 17, 9, 18, 19, 20, 21, 22]. As a consequence, PG methods have gained attention in recent studies focusing on the LQR problem. Research on these methods increased significantly after the work in [9] demonstrated global convergence of PG methods applied to the deterministic LQR problem. The work [13] explores a primal-dual policy gradient method to solve the constrained LQR problem. In [14], the performance limitations of PG applied to LQR problem are studied from a control-theoretic perspective.
@@
Solving the LQR problem requires the availability of a model of the system, as it is assumed in [11, 12, 13, 14]. However, in real-world applications, a complete system description is often unavailable, making it necessary to combine policy optimization schemes with data-driven techniques. To address this challenge, various approaches have been developed. For instance, methods like those in [23, 24, 25] combine system identification with model-based LQR design and use regret as a performance metric of the learning process. It is also possible to use the same two-step approach to solve the LQR problem with gradient-based schemes. The work [26] integrates recursive least squares with policy iteration optimization, while [27] combines recursive least squares with policy gradient methods. Both works [26, 27] assume noise-free data, whereas [28, 29] extend these frameworks to bounded-noise settings. These approaches are termed indirect data-driven methods, because first data is used to estimate a model, which is then integrated into a model-based certainty equivalence design.
Direct data-driven methods instead use data directly to determine the policy, by-passing the intermediate model estimation process [30]. In [9], model-free PG methods are proposed using zeroth-order optimization to estimate gradients for the LQR cost from finite but noise-free system trajectories. The work [31] proposed a novel gradient estimation method for noise-free systems. The work [32] uses PG methods to solve the continuous-time state feedback LQR problem. Similarly, in [33], PG methods are applied to a continuous-time output feedback system using zeroth-order gradient estimation alongside variance reduction techniques. The work [34] focuses on stabilizing the unknown system using policy gradient methods. In the aforementioned direct data-driven works [9, 31, 32, 33, 34], only the case where the system’s trajectories are noise-free is studied. However, it is essential to analyze what happens when gradients and other objects involved in the policy update are estimated using noisy data. In [10], model-free PG methods are applied to linear systems with multiplicative noise, addressing uncertainty in the system matrix. The works [10, 33, 34] all build on [9], leveraging the inherent robustness of PG methods: informally, when the gradient estimation error is sufficiently small, one can still guarantee a decrease in the cost function. However, in all those works [9, 10, 33, 34], stochastic uncertainty is not considered, and the analysis assumes that the trajectory data is strictly bounded. Another line of research tackles the model-free LQR problem differently. The work [35] derives convergence properties by carefully selecting step sizes when facing uncertainty of the gradient. Follow-up works [36, 37] refine the analysis by incorporating alternative gradient estimation techniques, yielding improved sample efficiency and robustness guarantees. These results [35, 36, 37] assume stochastic but bounded noise. There are also studies considering stochastic unbounded noise. For example, [38] applies PG methods to finite-horizon LQR, while [39, 40] propose model-free first-order methods, which estimate individual components of the gradient using least squares and differ from the aforementioned zeroth-order approaches. To date, there is no comprehensive analysis that jointly addresses the uncertainty introduced by the estimation process using the zeroth-order method and its propagation through the algorithm, and provides probabilistic convergence guarantees and sample complexity bounds for the infinite-horizon LQR problem.
@@
In this work, we analyze the sample complexity and robustness to noise of PG methods for the model-free LQR with additive stochastic noise. Due to the presence of noise with unbounded support, we consider an average infinite-horizon cost. We first characterize the properties of this cost function and analyze model-based versions of a few PG algorithms. For the policy gradient descent (PGD) and natural policy gradient (NPG) methods, we propose an adaptive stepsize scheme that has a provable faster convergence rate compared to the fixed step size methods discussed in [9, 10, 33]. Then, for both the model-free policy gradient descent and natural policy gradient methods, we employ zeroth-order optimization to approximate the gradient and covariance from noisy data. To handle the challenge of unbounded gradient estimates due to stochastic noise, we introduce an event-based analysis, extending the approach of [38], while relaxing the reliance on gradient upper bounds. We provide convergence guarantees for these model-free PG methods with a sample complexity that takes into account the noisy source of estimation error. Furthermore, we introduce and analyze a variance reduction technique within the zeroth-order optimization framework, and prove that this achieves a guaranteed improvement in the sample complexity. Even though variance reduction techniques for model-free LQR were already presented in [33], what we propose here is a new contribution compared to prior works, as we deal with stochastic data and we are able to provide a provable improvement in the number of required samples. Finally, we compare the guarantees of model-free PG methods with and without noisy data. This demonstrates the importance of considering noise, offers guidance on parameter tuning in applications, and also illustrates the improvement of introducing adaptive step sizes and variance reduction. Our analysis shows that model-free PG with stochastic data enjoys qualitatively similar convergence guarantees to those in the noise-free scenario, but it points out limitations in terms of the additional number of samples required for given accuracy, reduced step size ranges, and convergence rate.
@@
We denote by $A\succeq 0$ and $A\succ 0$ a positive semidefinite and positive definite matrix $A$, respectively. The symbol $\lambda_{1}(A)$ denotes the smallest eigenvalue of the matrix $A$. For matrices, $\lVert\cdot\rVert_{F}$ and $\lVert\cdot\rVert$ denote the Frobenius norm and induced $2$-norm, respectively. For vectors, $\lVert\cdot\rVert$ denotes the Euclidean norm. $I$ represents the identity matrix. $\mathbb{Z}_{+}$ is the set of non-negative integers. The operation $vec(A)=[a_{1}^{\top},a_{2}^{\top},...,a_{n}^{\top}]^{\top}$ stacks the columns of matrix $A$ into a vector. $\mathcal{N}(0,\Sigma)$ denotes a Gaussian distribution with 0 mean and covariance $\Sigma\succ 0$.
@@ 2 Preliminaries
$\min_{u_{t}}$ $\lim_{T\rightarrow+\infty}\frac{1}{T}\mathop{\mathbb{E}}_{x_{0},w_{t}}\sum_{t=0}^{T-1}\left(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}\right),$ (1a)
$\mathrm{s.\,t.}$~ $x_{t+1}=Ax_{t}+Bu_{t}+w_{t},w_{t}\sim\mathcal{N}(0,\Sigma_{w}),x_{0}\sim\mathcal{D},$ (1b)
where $x_{t}\in\mathbb{R}^{n_{x}}$ is the system state and $u_{t}\in\mathbb{R}^{n_{u}}$ is the system input; $A\in\mathbb{R}^{n_{x}\times n_{x}}$, $B\in\mathbb{R}^{n_{x}\times n_{u}}$, $(A,B)$ is unknown but stabilizable, which is a standard assumption in data-driven control [41, 42]; $Q,R\succ 0$ are the weighting matrix. We define the state covariance at the initial time as $\Sigma_{0}:=\mathop{\mathbb{E}}\limits_{x_{0}}\left[x_{0}x_{0}^{\top}\right]$ and then $\mathcal{D}:=\mathcal{N}(0,\Sigma_{0})$.
@@
$C(K):=\lim_{T\rightarrow+\infty}\frac{1}{T}\mathop{\mathbb{E}}_{x_{0},w_{t}}\sum_{t=0}^{T-1}x_{t}^{\top}\underbrace{\left(Q+K^{\top}RK\right)}_{=:Q_{K}}x_{t}.$ (2)
Remark 1 (Weighting matrices $Q,R$)
Prior knowledge of the weighting matrices $Q$ and $R$ is not strictly required for implementing the model-free policy gradient algorithms investigated here. While the analyses presented here assume their knowledge, it suffices to have access to the empirical cost, which can be estimated from the closed-loop trajectories of system (1b) used to estimate gradient information.
@@ 2.1 Properties of $C$
$\mathcal{S}:=\left\{K\in\mathbb{R}^{n_{x}\times n_{u}}|\lVert A_{K}\rVert<1\right\}.$ (4)
@@
$\Sigma_{t+1}=A_{K}\Sigma_{t}A_{K}^{\top}+\Sigma_{w},\quad t\in\mathbb{Z}_{+}.\\$ (5)
@@
$\Sigma_{K}:=\lim_{T\rightarrow+\infty}\frac{1}{T}\sum_{t=0}^{T-1}\Sigma_{t}.$ (6)
@@
$\Sigma_{K}=\Sigma_{w}+A_{K}\Sigma_{K}A_{K}^{\top}=\sum_{t=0}^{\infty}A_{K}^{t}\Sigma_{w}A_{K}^{t~\top}.$ (7)
@@
$C(K)=\mathrm{Tr}\left(P_{K}\Sigma_{w}\right)=\mathrm{Tr}\left(Q_{K}\Sigma_{K}\right),$ (8)
@@
It is known that the cost function $C$ defined in (2) is generally non-convex [9]. In the context of non-convex optimization, the convergence to the optimal solution of gradient descent cannot usually be guaranteed. However, the cost function $C(K)$ satisfies a special property known as gradient domination.
Lemma 1 (Gradient Domination)
The function $C$ on the set $\mathcal{S}$ is gradient dominated. That is, for any $K\in\mathcal{S}$, the following inequality holds:
$C(K)-C(K^{*})\leq\mu\lVert\nabla C(K)\rVert_{F}^{2},$ (12)
with $\mu:=\frac{1}{4}\lVert\Sigma_{K^{*}}\rVert\lVert\Sigma_{w}^{-2}\rVert\lVert R^{-1}\rVert$.
From (12) and the fact that $\Sigma_{w}\succ 0$, it follows that the cost function $C$ has a unique minimizer and no other stationary points. The proof of Lemma 1 is provided in Appendix A.1, and is a straightforward extension of [9, Lemma 3] to the averaged infinite horizon setting considered here (2).
Lemma 2 (Almost Smoothness on $\mathcal{S}$)
For any $K,K^{\prime}\in\mathbb{S}$, the following inequality holds:
$\begin{split}\left\lvert C(K^{\prime})-C(K)-2\mathrm{Tr}\left(({K^{\prime}}-{K})^{\top}E_{K}\Sigma_{K^{\prime}}\right)\right\rvert\leq\lVert\Sigma_{K^{\prime}}\rVert\left\lVert R+B^{\top}P_{K}B\right\rVert\left\lVert K^{\prime}-{K}\right\rVert_{F}^{2},\end{split}$ (13)
The proof of Lemma 2 is provided in Appendix A.2, which is an extension of [9, Lemma 6]. To understand why (13) is referred to as almost smoothness, assume additionally that the term $\lVert\Sigma_{K}\rVert,\lVert R+B^{\top}P_{K}B\rVert$ can be upper bounded by a constant $L$, and that $K^{\prime}$ is sufficiently close to $K$ so that we can approximate $\Sigma_{K}$ with $\Sigma_{K^{\prime}}$, i.e., $\Sigma_{K}\approx\Sigma_{K^{\prime}}$. Then, recalling (13), we see that (13) is equivalent to:
@@ 2.2 Perturbation of $\Sigma_{K},C,\nabla C$
In this subsection, we demonstrate that the average covariance matrix $\Sigma_{K}$, the cost function $C(K)$, and its gradient $\nabla C(K)$ are locally Lipschitz continuous with respect to the policy $K$. This property is crucial for the analysis of model-free policy optimization.
Lemma 3 ($\Sigma_{K}$ Perturbation)
Suppose $K^{\prime},K\in\mathcal{S}$ are such that:
@@
The proof of Lemma 3 is given in Appendix A.3, which follows the proof of [9, Lemma 16] and uses the definition of $\Sigma_{K}$.
Lemma 4 ($C$ Perturbation)
Suppose $K^{\prime},K\in\mathcal{S}$ are such that:
@@
where $h_{C}(C(K))$ is defined in (55) in the proof.
Lemma 5 ($\nabla C$ Perturbation)
Suppose $K^{\prime}$ is such that:
@@
where $h_{\nabla}(C(K))$ is defined in (59) in the proof.
In summary, Lemma 3, Lemma 4 and Lemma 5 establish that $\Sigma_{K}$, $C$ and $\nabla C$ are locally Lipschitz continuous. The proofs of Lemma 4 and Lemma 5 are provided in Appendix A.4 and A.5, respectively. These proofs follow the approach used in [9, Lemma 24, Lemma 25] by adapting the expressions of $C$ and $\nabla C$.